OpenStudy (anonymous):

steps to find the nth derivative of y = cos5x.cos3x

OpenStudy (chaise):

Is that dot used to represent multiplication? \[y=\cos5x \times \cos3x\]

OpenStudy (anonymous):

f(x) = g(x) y(x) f'(x) = g(x)y'(x) + f(x) g'(x)

sam (.sam.):

They're steps too

OpenStudy (anonymous):

mul wiht 2and divide with 2 \[{\cos(5x+3x)+\cos(5x-2x)}/2\] or \[(\cos7x+\cos3x)/2\] then apply nth derivative